# Collection: Datasets from AFAR Challenge

### v1

@@ Abstract
This paper presents a comprehensive real-world and Digital Twin (DT) dataset collected as part of the Find A Rover (AFAR) Challenge, organized by the NSF Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) testbed and hosted at the Lake Wheeler Field in Raleigh, North Carolina. The AFAR Challenge was a competition involving five finalist university teams, focused on promoting innovation in UAV-assisted radio frequency (RF) source localization. Participating teams were tasked with designing UAV flight trajectories and localization algorithms to detect the position of a hidden unmanned ground vehicle (UGV), also referred to as a rover, emitting wireless probe signals generated by GNU Radio. The competition was structured to evaluate solutions in a DT environment first, followed by deployment and testing in AERPAW’s outdoor wireless testbed. For each team, the UGV was placed at three different positions, resulting in a total of 30 datasets—15 collected in a DT simulation environment and 15 in a physical outdoor testbed. Each dataset contains time-synchronized measurements of received signal strength (RSS), received signal quality (RSQ), GPS coordinates, UAV velocity, and UAV orientation (roll, pitch, and yaw). Data is organized into structured folders by team, environment (DT and real-world), and UGV location. The dataset supports research in UAV-assisted RF source localization, air-to-ground (A2G) wireless propagation modeling, trajectory optimization, signal prediction, autonomous navigation, and DT validation. With approximately 300k time-synchronized samples collected from real-world experiments, the dataset provides a substantial foundation for training and evaluating deep learning (DL) models. Overall, the AFAR dataset serves as a valuable resource for advancing robust, real-world solutions in UAV-enabled wireless communications and sensing systems.
[…]
DATA TYPE/LOCATION Time-series; AERPAW Lake Wheeler Field Labs, Raleigh, North Carolina, USA; AERPAW Digital Twin for Lake Wheeler Field Labs
{IEEEkeywords}
Air-to-ground propagation modeling, Digital twin, Real-world wireless datasets, UAV-assisted localization, UAV communication systems
@@
Unmanned aerial vehicles (UAVs) are integral to location-based applications such as search and rescue, surveillance, and radio frequency (RF) source localization due to their autonomous 3D navigation, high-altitude signal reception, and superior mobility compared to terrestrial networks. However, most research involving UAVs for assisting rescue operations, localization, or advancing wireless communication is conducted in simulated environments that do not fully capture real-world complexities, leading to challenges in translating solutions to practical deployments. Accurate algorithm development for UAV-assisted applications must account for real-world channel dynamics, mobility variations, and environmental effects—factors often overlooked in simulation-based studies. However, the lack of testbed facilities limits comprehensive validation and testing in a real-world environment. To bridge this gap, there is a growing need for integrated platforms that support both virtual development environments and real-world testing. These platforms enable researchers to design, implement, and iteratively refine solutions in controlled digital settings before transitioning to field deployment. The Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) addresses this need by offering a unique dual-environment testbed—comprising a Digital Twin (DT) framework for virtual development and experimentation, and a physical outdoor wireless testbed for real-world execution. As illustrated in Fig. 2, AERPAW facilitates seamless transition from simulation to deployment, supporting scalable and repeatable evaluation of UAV-assisted wireless systems.
@@
0.1 AFAR Competition
AERPAW hosted the AERPAW Find A Rover (AFAR) Challenge, inviting students from universities across the United States to bridge the gap between research and practical implementation in UAV-assisted wireless localization. The challenge tasked participants with designing and developing innovative algorithms to localize an unmanned ground vehicle (UGV), also referred to as a rover, using a UAV-based receiver. The UAV, acting as the signal receiver, collects RF signals during flight, while the UGV serves as the RF source transmitter. The resulting dataset includes time-synchronized measurements such as RSS, RSQ, GPS coordinates, and UAV orientation data, making it a valuable resource for spatiotemporal analysis and algorithm validation.
The experimental setup allowed the UGV to be placed anywhere within a designated area (marked in green in Fig. 2), while the UAV was restricted to flying within a predefined flight zone (marked in blue). A sample UAV trajectory is overlaid within the flight zone. Notably, portions of the green area fall outside the blue region, forming designated no-fly zones. These constraints prevent the UAV from flying directly above certain UGV positions, thereby making the localization problem more challenging, necessitating advanced localization strategies along with optimized trajectory planning. Teams were allowed to develop either an autonomous flight trajectory or a predefined waypoint-based trajectory for the UAV. The UAV was permitted to fly at altitudes ranging from 20 m to 110 m, with a maximum speed of 10 m/s.
Figure 2: Experimental setup of the AFAR Challenge, illustrating the UAV flight zone (blue) and the region where the UGV can be positioned (green), with the UAV utilized in the experiment depicted in the top right corner.
0.2 Scoring and Ranking
@@
0.3 Dataset Overview and Uniqueness
The AFAR Challenge dataset contains data collected from the top five teams that participated in the competition, with each team developing distinct UAV flight trajectories and localization algorithms to estimate the position of a UGV. The dataset includes both real-world and DT data, consistently structured across both environments. For each team, the UGV was deployed at three different locations (Loc-1, Loc-2, Loc-3), resulting in a total of 30 datasets—15 from real-world experiments and 15 from the DT environment. The participating teams were Team SunLab from the University of Georgia (T-288), Team NYU Wireless from New York University (T-300), Team Eagles from the University of North Texas (T-301), Team Wolfpack from NC State University (T-309), and Team Daedalic Wings from NC State University (T-328).
To the best of the authors’ knowledge, no prior challenge of this nature has been organized, nor has such a dataset been publicly released. This initiative represents the first large-scale dataset designed to facilitate advanced wireless communication analysis using UAVs, with a particular focus on the RF localization and air-to-ground (A2G) channel characterization. While RF source localization has been explored in previous studies, it has predominantly been investigated in simulated environments [2, 3, 4]. Some experimental efforts have focused on time difference of arrival-based localization techniques as in [5], which examines the impact of altitude, bandwidth, and NLOS bias on 3D UAV localization. Meanwhile, the studies in [6] and [7] examine A2G channel characteristics through experimental measurements conducted in a rural environment, where recorded signal data was later used to refine path loss models. However, this dataset is not publicly available, and the study primarily considers a static RF source at a fixed location, focusing on channel characterization rather than localization.
0.4 How the Dataset can be Reused
This dataset plays a pivotal role in advancing A2G channel modeling, capturing key characteristics such as propagation effects, fading impact, noise variations, and the effects of UAV’s speed and orientation. The data provides critical insights into UAV-assisted wireless communications, making it a valuable resource for multiple research domains. The key contributions of the AFAR dataset are as follows.
@@ 0.4 How the Dataset can be Reused
The dataset collected during the challenge was utilized by participating teams to develop and evaluate their localization algorithms. The AFAR dataset enabled teams to assess the performance of their approaches under practical conditions. A detailed analysis of the localization methodologies employed by the top three teams, along with their utilization of the AFAR dataset, is provided in [8], which also offers an overview of the challenge framework. Additionally, another participating team documented their approach in [9], while the implementation details of the third-ranking team are available in [10]. The study in [11] leveraged the AFAR dataset to bridge the gap between simulation and real-world deployment. The authors proposed an enhanced two-ray path loss model to improve A2G channel modeling by incorporating shadowing effects induced by the UAV body. To accurately capture these effects, they utilized the roll, yaw, and pitch angles provided in the dataset. The AFAR dataset was also used to validate a DL-based localization model trained on simulated data, demonstrating its effectiveness in real-world conditions.
@@
Data collection for the AERPAW AFAR challenge begins with team members signing up on the AERPAW experimenter web portal [12]. After registering, they join a project created by their project principal investigator (PI). Within the platform’s DT environment, each team initiates its experiment by choosing two nodes: one UGV, also known as the rover, and one UAV. In this setup, the UGV runs the transmitter part of the AERPAW channel sounder sample experiment [13], while the UAV runs the receiver part. The goal is to find the location of the rover based on the channel sounder measurements at the receiver. Instructions for manually moving a vehicle in AERPAW are used to “hide” the rover. UAV does not know the location of the rover, and it tries to detect its position based solely on the RSS measurements at the UAV.
Participants get themselves familiar with the AFAR challenge by first trying the AERPAW to Find a Rover Sample Code [13]. In the sample code, the UAV moves forward as long as the RSS improves; if the signal strength decreases, it turns 90 degrees to the right. Although this simple algorithm does not guarantee accurate location of the rover, it introduces participants to key concepts: how to read RSS measurements and how to control the movement of the UAV within geofence restrictions imposed by AERPAW. As the UAV approaches the rover, RSS generally increases. However, due to the nature of wireless channel variations and the influence of antenna patterns, the RSS does not always increase monotonically. This variation in signal strength poses a challenge in designing robust localization algorithms. Participants develop their strategies in the AERPAW DT—an environment that emulates the behavior of the real-world testbed. While the DT’s channel characteristics do not exactly mirror reality, AERPAW supplements this limitation by providing real testbed data. These datasets include RSS as a function of distance between the transmitter and the receiver, giving teams valuable insight into real-world propagation behavior. Once all teams finalize their algorithms, the code is first tested in the DT, and the same code (in software containers) is moved as it is to the physical testbed. The evaluation in the DT also helped AERPAW operators to identify and eliminate potential bugs before running the experiment in the testbed.
@@
Table 1: Statistics of RSS, signal RSQ, and UAV speed collected at Loc-1, Loc-2, and Loc-3 for team-288.
@@
The AFAR dataset exhibits high completeness and consistency, with no missing entries or out-of-range values across its features. All recorded parameters remain within expected operational bounds. Table 1 presents the statistical characteristics of RSS, RSQ, and UAV speed for Team 288 across the three deployment locations. The observed values are well aligned with the expected behavior of UAV-assisted RF measurements. Specifically, both RSS and RSQ demonstrate a wide dynamic range, reflecting realistic variations caused by the UAV’s relative proximity to the RF source. Stronger signal values are recorded when the UAV is closer to the RF transmitter, whereas lower values correspond to increased separation distances. The UAV speed during the flights ranges from approximately 0.001 m/s to 5 m/s. This variation is attributed to the waypoint-based trajectory design adopted by the teams. As the UAV reaches a waypoint, it momentarily decelerates or halts, resulting in near-zero speed readings. Subsequently, as it transitions to the next waypoint, the UAV accelerates, producing higher speed values. All other features (further detailed in Section IV) are similarly well-bounded and free from anomalies.
Table 2: Statistics of received samples.
@@
Throughout all experiments, RSS and RSQ measurements were consistently sampled at a fixed rate of 30 samples per second. Table 2 provides statistics on the sampling intervals and completeness of the collected data. It reports the total flight duration for each team and location, the expected number of samples based on the nominal sampling rate, and the actual number of samples recorded. In some instances, the number of collected samples slightly exceeds the expected count due to the equipment operating at a marginally higher sampling rate. Conversely, minor deficits in a few cases are observed but remain negligible, demonstrating the precision, reliability, and completeness of the data collection process. This consistency ensures the dataset’s suitability for accurate and reproducible analysis. For Team 309 at Location 1, the dataset is unavailable due to accidental deletion during file transfer. While the localization estimate is still available, the raw data was lost due to a handling error, not related to any issue in data collection or recording.
The UAV state features, including longitude, latitude, altitude, and speed, were sampled at a lower rate of approximately 5 Hz, whereas the RSS and RSQ measurements were acquired at a higher rate of 30 Hz. This sampling disparity introduces the challenge of associating each RSS and RSQ sample with the corresponding spatial and kinematic information from the UAV. To address this, linear interpolation was employed to estimate the UAV state feature at the higher sampling rate of the RSS and RSQ measurements. For example, for each RSS sample timestamp, the corresponding UAV altitude was computed by interpolating between the available GPS altitude samples. This interpolation process was applied similarly to other UAV state features, ensuring temporal alignment across all recorded features.
Fig. 3 illustrates the result of the interpolation process, where the original sparse GPS altitude samples (in blue) are complemented by interpolated values (in red). The interpolated points accurately follow the trend of the original GPS measurements without introducing noticeable bias or drift, thereby preserving the physical consistency of the UAV trajectory. This interpolation strategy enables the seamless integration of all features during post-processing and ensures that each RSS and RSQ measurement is reliably associated with the corresponding information.
The GPS-derived coordinates of the UAV, including longitude, latitude, and altitude, are observed to be within the predefined flight boundary, as expected. Fig. 4 illustrates the UAV trajectory for Team-288 for the UGV located at Location-1. It can be clearly seen that all recorded UAV positions are confined within the designated bounding box, confirming the correctness of the recorded navigation data. The plotted trajectory (shown in blue) effectively traces the UAV’s movement, while the RF source location is marked and remains stationary. This visualization also highlights the UAV’s maneuvering around the RF source, validating the integrity of the positional data.
@@
Table 3: Description of primary data files in each location-specific folder.
@@
$power\_log.txt$ Records time-synchronized RSS values. Contains three columns: the first for timestamps and the third for RSS.
$quality\_log.txt$ Captures RSQ values with timestamps. Structured similarly to the power log.
@@
While the primary focus of this paper is on the analysis and exploration of the real-world (testbed) dataset, both the emulation (DT) and testbed datasets share a unified structure. The dataset is organized into directories based on each team’s experiment identification number (T-XXX). Within each team folder, there are two subfolders: ‘development’, containing the DT-collected data, and ‘testbed ’, containing the real-world data. Each of these subfolders is further divided into three location-specific directories—loc-1, loc-2, and loc-3—corresponding to different UGV placements (e.g., 288/loc-1, 288/loc-2, 288/loc-3). For the real-world dataset located in the ‘testbed’ folder, each location-specific directory contains four primary files: ‘$log.csv$’, ‘$power\_log.txt$’, ‘$quality\_log.txt$‘, and ‘$angles.mat$’, as summarized in Table 3. These files include timestamped measurements, enabling chronological alignment of sensor data and supporting detailed analyses of channel coherence time, UAV-assisted RF localization, and A2G propagation modeling, particularly under varying mobility patterns, environmental conditions, and UAV orientations.
Table 4 provides a detailed description of the fields contained in the $log.csv$ file, which records the UAV’s navigation and positional data throughout each flight. These fields include highly time-synchronized GPS-derived coordinates (latitude, longitude, and altitude), speed metrics, and satellite information. The combination of temporal and spatial parameters allows for the accurate reconstruction of the UAV’s flight trajectory and enables synchronization with RF measurements such as RSS and RSQ. Table 5 outlines the structure of the power_log.txt file, which contains RSS measurements captured by the UAV during flight. The quality_log.txt file follows an identical format, with the third column containing RSQ values instead of RSS. Together, these files offer a comprehensive dataset for characterizing and analyzing A2G wireless channel performance.
Table 4: Description of key fields in $log.csv$, capturing UAV state features.
@@
Table 5: Description of fields in power_log.txt file.
@@
Table 6 summarizes the UAV orientation parameters available in the angles.mat file. These values enable precise attitude estimation, which is essential for modeling the shadowing effects and understanding the impact of UAV dynamics on signal propagation and localization performance.
Table 6: Description of variables in angles.mat file.
@@
The AFAR dataset is a highly versatile and comprehensive resource for advancing UAV-assisted wireless communication research. It encompasses diverse UAV flight trajectories with UGV at multiple locations, providing rich spatial and temporal variations. The dataset includes all essential features required for A2G propagation modeling and real-world testing of RF localization algorithms under practical environmental conditions. By capturing both DT simulations and real-world measurements, the dataset offers valuable insight into the discrepancies between simulated and actual deployment scenarios. This duality allows researchers to quantify and address the simulation-to-reality gap, enabling the development of more robust and adaptive models.
@@
An illustrative example of the AFAR dataset’s richness and utility is presented in Fig. 5, where multiple features—including UAV speed, altitude, trajectory, distance to the RF source, and RSS—are visualized for all participating teams at Loc-2. This figure offers detailed insights into the dataset’s structure and demonstrates how it can be leveraged to explore the relation between key variables. The first column of the figure displays UAV altitude and speed as functions of time. The second column presents a heatmap of RSS values overlaid on the UAV’s flight trajectory, illustrating spatial variations in RSS across the area of operation. The third column compares real-world and DT RSS measurements over time, alongside the temporal variation in distance between the UAV and the RF source.
@@
From the first column in Fig. 5, it can be observed that each team employed different altitudes for UAV flight during their respective trajectories. In all cases, the UAV ascends from its starting position at the beginning of the flight and descends near the end, indicating the takeoff and landing phases. Outside of these transitions, the UAV typically maintains a constant altitude. Altitude plays a significant role in RSS behavior—higher altitudes improve line-of-sight (LOS) conditions; however, greater altitude also increases the distance to the RF source, which can lead to signal attenuation. The AFAR dataset allows analysis of RSS variation with UAV altitude, providing real-world insights.
@@
The same column also presents the UAV’s speed profile, which exhibits a consistent pattern of increasing and decreasing speed over time. The team’s UAV exhibits a speed range from 0 to 10 m/s, reflecting diverse operational states and maneuvering behaviors throughout the flight. This is attributed to the waypoint-based UAV trajectory. UAV trajectories in AFAR are of two types: fixed waypoints and autonomous waypoints. In the fixed approach, teams predefine a set of waypoints in a plan file that the UAV follows sequentially. In contrast, the autonomous approach involves decision-making algorithms that dynamically determine the next waypoint based on the current state and measurements. These two strategies lead to distinct movement patterns and influence the UAV’s interaction with the RF environment, as can be seen in the second column of Fig. 5. The UAV accelerates while traveling toward a waypoint and decelerates upon arrival, where it may pause briefly before proceeding, causing a decrease in the speed, which is visible in Fig. 5. This behavior is consistent across teams and trajectories. UAV speed also has an important impact on signal measurement. Variations in speed can affect RSS sampling consistency and localization accuracy. Using the AFAR dataset, researchers can investigate the relationship between UAV speed, RSS fluctuations, and overall system performance, enabling more informed and adaptive algorithm development.

### v2

# Collection: UAV-Based RSS Measurements from the AFAR Challenge in Digital Twin and Real-World Environments

@@ Abstract
This paper presents a comprehensive real-world and Digital Twin (DT) dataset collected as part of the AERPAW Find A Rover (AFAR) Challenge, organized by the NSF Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) testbed and hosted at the Lake Wheeler Field in Raleigh, North Carolina. The AFAR Challenge was a competition involving five finalist university teams, focused on promoting innovation in unmanned aerial vehicle (UAV)-assisted radio frequency (RF) source localization. Participating teams were tasked with designing UAV flight trajectories and localization algorithms to detect the position of a hidden unmanned ground vehicle (UGV), also referred to as a rover, emitting probe signals generated by GNU Radio. The competition was structured to evaluate solutions in a DT environment first, followed by deployment and testing in AERPAW's outdoor wireless testbed. For each team, the UGV was placed at three different positions, resulting in a total of 29 datasets—15 collected in a DT simulation environment and 14 in a physical outdoor testbed. Each dataset contains time-synchronized measurements of received signal strength (RSS), received signal quality (RSQ), GPS coordinates, UAV velocity, and UAV orientation (roll, pitch, and yaw). Data is organized into structured folders by team, environment (DT and real-world), and UGV location. The dataset supports research in UAV-assisted RF source localization, air-to-ground (A2G) wireless propagation modeling, trajectory optimization, signal prediction, autonomous navigation, and DT validation. With 300k time-synchronized samples from the real-world experiments, the AFAR dataset enables effective training/testing of deep learning (DL) models and supports robust, real-world UAV-based wireless communication and sensing research.
IEEE SOCIETY/COUNCIL: IEEE Communications Society (ComSoc), IEEE Vehicular Technology Society (VTS), IEEE Aerospace and Electronic Systems Society (AESS)
DATA DOI/PID: 10.5061/dryad.18931zd4g
DATA TYPE/LOCATION: Time-series; AERPAW Lake Wheeler Field Labs, Raleigh, North Carolina, USA; AERPAW Digital Twin for Lake Wheeler Field Labs
@@
Unmanned aerial vehicles (UAVs) are integral to location-based applications such as search and rescue, surveillance, and radio frequency (RF) source localization due to their autonomous 3D navigation, high-altitude signal reception, and superior mobility compared to terrestrial networks. However, most research involving UAVs for assisting rescue operations, localization, or advancing wireless communication is conducted in simulated environments that do not fully capture real-world complexities, leading to challenges in translating solutions to practical deployments. Accurate algorithm development for UAV-assisted applications must account for real-world channel dynamics, mobility variations, and environmental effects—factors often overlooked in simulation-based studies. However, the lack of testbed facilities limits comprehensive validation and testing in a real-world environment. To bridge this gap, there is a growing need for integrated platforms that support both virtual development environments and real-world testing. These platforms enable researchers to design, implement, and iteratively refine solutions in controlled digital settings before transitioning to field deployment. The Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) addresses this need by offering a digital twin (DT) for virtual experimentation and a physical outdoor testbed for real-world deployment. As illustrated in Fig. 1, AERPAW facilitates seamless transition from simulation to deployment, supporting scalable and repeatable evaluation of UAV-assisted wireless systems. The aforementioned figure is adapted with modifications from [1] to highlight the differences between the DT and real-world environments.
@@
AFAR Competition
AERPAW hosted the AERPAW Find A Rover (AFAR) Challenge, inviting students from universities across the United States to bridge the gap between research and practical implementation in UAV-assisted RF source localization. The challenge tasked participants with designing and developing innovative algorithms to localize an unmanned ground vehicle (UGV), also referred to as a rover, using a UAV-based receiver. The UAV, a custom hexacopter shown in the upper-right corner of Fig. 2, collects RF signals during flight, while the UGV serves as the RF source transmitter. The resulting dataset includes time-synchronized measurements such as RSS, RSQ, GPS coordinates, and UAV orientation data, making it a valuable resource for spatiotemporal analysis and algorithm validation.
The experimental setup allowed the UGV to be placed anywhere within a designated area (marked in green in Fig. 2), while the UAV was restricted to flying within a predefined flight zone (marked in blue). A sample UAV trajectory is overlaid within the flight zone. Notably, portions of the green area fall outside the blue region, forming designated no-fly zones. These constraints prevent the UAV from flying directly above certain UGV positions, thereby making the localization problem more challenging, necessitating advanced localization strategies along with optimized trajectory planning. Teams were allowed to develop either an autonomous flight trajectory or a predefined waypoint-based trajectory for the UAV. The UAV was permitted to fly at altitudes ranging from 20 m to 110 m, with a maximum speed of 10 m/s. The real-world measurements were collected at the Lake Wheeler Field, a rural outdoor test site that includes large open grass fields, surrounding tree lines, scattered bushes, and limited man-made structures, as shown in Fig. 2.
Figure 2: Experimental setup of the AFAR Challenge, illustrating the UAV flight zone (blue) and the region where the UGV can be positioned (green), with the UAV used in the experiment depicted in the top right corner.
Scoring and Ranking
@@
Dataset Overview and Uniqueness
The AFAR Challenge dataset contains data collected from the top five teams that participated in the competition, with each team developing distinct UAV flight trajectories and localization algorithms to estimate the position of a UGV. The dataset includes both real-world and DT data, consistently structured across both environments. For each team, the UGV was deployed at three different locations (Loc-1, Loc-2, Loc-3), resulting in a total of 30 datasets, with 14 from real-world experiments and the remaining 15 from the DT environment. The participating teams were Team SunLab from the University of Georgia (T-288), Team NYU Wireless from New York University (T-300), Team Eagles from the University of North Texas (T-301), Team Wolfpack from NC State University (T-309), and Team Daedalic Wings from NC State University (T-328).
To the best of the authors’ knowledge, no prior challenge of this nature has been organized, nor has such a dataset been publicly released. This initiative represents the first large-scale dataset designed to facilitate advanced wireless communication analysis using UAVs, with a particular focus on the RF localization and air-to-ground (A2G) channel characterization. While RF source localization has been explored in previous studies, it has predominantly been investigated in simulated environments [2, 3, 4]. Some experimental efforts have focused on time difference of arrival-based localization techniques, as in [5], which examines the impact of altitude, bandwidth, and non-line-of-sight (NLOS) bias on 3D UAV localization. Meanwhile, the studies in [6] and [7] examine A2G channel characteristics through experimental measurements conducted in a rural environment, where recorded signal data was later used to refine path loss models. However, this dataset [6] is not publicly available, and the study primarily considers a single, fixed-location RF source for channel characterization rather than localization. In contrast, our dataset uses a stationary RF source placed at multiple distinct locations across trials, enabling analysis and benchmarking of UAV-based RF localization algorithms.
How the Dataset can be Reused
This dataset plays a pivotal role in advancing UAV-assisted RF source localization by enabling researchers to design, train, and benchmark algorithms under diverse flight paths and propagation conditions. In addition, it supports A2G channel modeling by capturing key characteristics such as propagation effects, fading, noise variations, and the influence of UAV speed and orientation. Together, these features provide critical insights into UAV-assisted wireless communication systems, making the dataset a valuable resource for a wide range of research domains. The key contributions of the AFAR dataset are as follows.
@@
The dataset collected during the challenge was utilized by participating teams to develop and evaluate their localization algorithms. The AFAR dataset enabled teams to assess the performance of their approaches under practical conditions. A detailed analysis of the localization methodologies employed by the top three teams, along with their utilization of the AFAR dataset, is provided in [8], which also offers an overview of the challenge framework. Additionally, another participating team documented their approach in [9], while the implementation details of the third-ranking team are available in [10]. The study in [11, 12] leveraged the AFAR dataset to bridge the gap between simulation and real-world deployment. The authors in [11] proposed an enhanced two-ray path loss model to improve A2G channel modeling by incorporating shadowing effects induced by the UAV body. To accurately capture these effects, they utilized the roll, yaw, and pitch angles provided in the dataset. The AFAR dataset was also used to validate a DL-based localization model trained on simulated data, demonstrating its effectiveness in real-world conditions.
@@
Within the platform’s DT environment, each team initiates its experiment by choosing two nodes: one UGV, also known as the rover, and one UAV. In this setup, the UGV runs the transmitter part of the AERPAW channel sounder sample experiment [14], while the UAV runs the receiver part. UAV does not know the location of the rover, and it tries to detect its position based solely on the RSS measurements at the UAV. As the UAV approaches the rover, RSS generally increases. However, due to the nature of wireless channel variations and the influence of antenna patterns, the RSS does not always increase monotonically. This variation in signal strength poses a challenge in designing robust localization algorithms. Participants develop their strategies in the AERPAW DT—an environment that emulates the behavior of the real-world testbed. While the DT’s channel characteristics do not exactly mirror reality, AERPAW supplements this limitation by providing real testbed data. These datasets include RSS as a function of distance between the transmitter and the receiver, giving teams valuable insight into real-world propagation behavior. Once all teams finalize their algorithms, the code is first tested in the DT, and the same code (in software containers) is moved as it is to the physical testbed. The evaluation in the DT also helped AERPAW operators to identify and eliminate potential bugs before running the experiment in the testbed.
@@
TABLE I: Statistics of RSS, signal RSQ, and UAV speed collected at Loc-1, Loc-2, and Loc-3 for team-288.
@@
The AFAR dataset exhibits high completeness and consistency, with no missing entries or out-of-range values across its features. All recorded parameters remain within expected operational bounds. Table I presents the statistical characteristics of RSS, RSQ, and UAV speed for Team 288 across the three deployment locations. These results are shown as a representative example; for the other teams, the mean and standard deviation values differ due to variations in their designed UAV trajectories as well as additional factors such as multipath reflections and noise. Specifically, both RSS and RSQ exhibit a wide dynamic range, which arises not only from changes in the UAV’s distance to the RF source but also from propagation effects such as multipath fading, environmental shadowing, noise, and UAV body or orientation-induced blockage. The UAV speed during the flights ranges from approximately 0.001 m/s to 5 m/s. This variation is attributed to the waypoint-based trajectory design adopted by the teams. As the UAV reaches a waypoint, it momentarily decelerates or halts, resulting in near-zero speed readings. Subsequently, as it transitions to the next waypoint, the UAV accelerates, producing higher speed values. All other features (further detailed in Section IV) are similarly well-bounded and free from anomalies.
TABLE II: Statistics of received samples.
@@
Throughout all experiments, RSS and RSQ measurements were consistently sampled at a fixed rate of 30 samples per second. Table II provides statistics on the sampling intervals and completeness of the collected data. It reports the total flight duration for each team and location, the expected number of samples based on the nominal sampling rate, and the actual number of samples recorded. In some instances, the number of collected samples slightly exceeds the expected count due to the equipment operating at a marginally higher sampling rate. Conversely, minor deficits in a few cases are observed but remain negligible, demonstrating the precision, reliability, and completeness of the data collection process. These variations do not affect time stamp accuracy, as they are system-clock-generated and aligned with actual sampling instances. For Team 309 at Location 1, the physical testbed dataset is unavailable due to accidental deletion during file transfer, although the localization estimate remains available. Locations 2 and 3 are intact, and the DT data for all three locations is fully available. Thus, only one physical dataset is missing, while overall coverage and reusability for benchmarking remain unaffected.
For Team 309 at Location 1, the dataset is unavailable due to accidental deletion during file transfer. While the localization estimate is still available, the raw data was lost due to a handling error, not related to any issue in data collection or recording.
@@
The UAV state features, including longitude, latitude, altitude, speed, and absolute heading, were sampled at a lower rate of approximately 5 Hz through the flight computer. Positioning and timing are provided by an onboard real-time kinematic (RTK)-enabled global navigation satellite system (GNSS) system, with centimeter-level accuracy ensured by RTK corrections from AERPAW’s base station at Lake Wheeler. In contrast, the RSS and RSQ measurements were acquired at a higher rate of 30 Hz. This sampling disparity introduces the challenge of associating each RSS and RSQ sample with the corresponding spatial and kinematic information from the UAV. To address this, linear interpolation was employed to estimate the UAV state feature at the higher sampling rate of the RSS and RSQ measurements. For example, for each RSS sample timestamp, the corresponding UAV altitude was computed by interpolating between the available GPS altitude samples. This interpolation process was applied similarly to other UAV state features, ensuring temporal alignment across all recorded features. Fig. 3 illustrates the result of the interpolation process, where the original sparse GPS altitude samples (in blue) are complemented by interpolated values (in red). The interpolated points closely follow the trend of the original GPS samples, ensuring consistency of the UAV trajectory and reliable alignment of RSS/RSQ with state features.
The GPS-derived coordinates of the UAV, including longitude, latitude, and altitude, are observed to be within the predefined flight boundary, as expected. Fig. 4 illustrates the UAV trajectory for Team-288 for the UGV located at Location-1. This visualization also highlights the UAV’s maneuvering around the RF source, validating the integrity of the positional data.
@@
TABLE III: Description of primary data files in each location-specific folder.
@@
power_log.txtpower\_log.txt Records time-synchronized RSS values. Contains three columns: the first for timestamps and the third for RSS.
quality_log.txtquality\_log.txt Captures RSQ values with timestamps. Structured similarly to the power log.
@@
While the primary focus of this paper is on the analysis and exploration of the real-world dataset, both the DT and real-world datasets share a unified structure. The dataset is organized into directories based on each team’s experiment ID (T-XXX). Within each team folder, there are two subfolders: ‘development’, containing the DT-collected data, and ‘testbed ’, containing the real-world data. Each of these subfolders is further divided into three location-specific directories—loc-1, loc-2, and loc-3, corresponding to different UGV placements (e.g., 288/loc-1, 288/loc-2, 288/loc-3). For the real-world dataset located in the ‘testbed’ folder, each location-specific directory contains four primary files: ‘$log.csv$’, ‘power_log.txtpower\_log.txt’, ‘quality_log.txtquality\_log.txt‘, and ‘$angles.mat$’, as summarized in Table III.
Table IV provides a detailed description of the fields contained in the $log.csv$ file, which records the UAV’s navigation and positional data throughout each flight. These fields include highly time-synchronized GPS-derived coordinates (latitude, longitude, and altitude), speed metrics, and satellite information. The combination of temporal and spatial parameters allows for the accurate reconstruction of the UAV’s flight trajectory and enables synchronization with RF measurements such as RSS and RSQ. Table V outlines the structure of the power_log.txt file, which contains RSS measurements captured by the UAV during flight. The quality_log.txt file follows an identical format, with the third column containing RSQ values instead of RSS. Together, these files offer a comprehensive dataset for characterizing and analyzing A2G wireless channel performance.
TABLE IV: Description of key fields in $log.csv$, capturing UAV state features.
@@
TABLE V: Description of fields in power_log.txt file.
@@
Table VI summarizes the UAV orientation parameters available in the angles.mat file. These values enable precise attitude estimation, which is essential for modeling the shadowing effects and understanding the impact of UAV dynamics on signal propagation and localization performance.
TABLE VI: Description of variables in angles.mat file.
@@
The AFAR dataset is a highly versatile and comprehensive resource for advancing UAV-assisted wireless communication research. An illustrative example of the AFAR dataset’s richness and utility is presented in Fig. 5, where multiple features—including UAV speed, altitude, trajectory, distance to the RF source, and RSS—are visualized for all participating teams at Loc-2. This figure offers detailed insights into the dataset’s structure and demonstrates how it can be leveraged to explore the relation between key variables. The first column of the figure displays UAV altitude and speed as functions of time. The second column presents a heatmap of RSS values overlaid on the UAV’s flight trajectory, illustrating spatial variations in RSS across the area of operation. The third column compares real-world and DT RSS measurements over time, alongside the temporal variation in distance between the UAV and the RF source.
@@
From the first column in Fig. 5, it can be observed that each team employed different altitudes for UAV flight during their respective trajectories. In all cases, the UAV ascends from its starting position at the beginning of the flight and descends near the end, indicating the takeoff and landing phases. Outside of these transitions, the UAV typically maintains a constant altitude. Altitude plays a significant role in RSS behavior—higher altitudes improve LoS conditions; however, greater altitude also increases the distance to the RF source, which can lead to signal attenuation. The AFAR dataset allows analysis of RSS variation with UAV altitude, providing real-world insights.
@@
The same column also presents the UAV’s speed profile, which exhibits a consistent pattern of increasing and decreasing speed over time. The team’s UAV exhibits a speed range from 0 to 10 m/s, reflecting diverse operational states and maneuvering behaviors throughout the flight. This is attributed to the waypoint-based UAV trajectory. UAV trajectories in AFAR are of two types: fixed waypoints and autonomous waypoints. In the fixed approach, teams predefine a set of waypoints in a plan file that the UAV follows sequentially. In contrast, the autonomous approach involves decision-making algorithms that dynamically determine the next waypoint based on the current state and measurements. These two strategies lead to distinct movement patterns and influence the UAV’s interaction with the RF environment, as can be seen in the second column of Fig. 5. The UAV accelerates while traveling toward a waypoint and decelerates upon arrival, where it may pause briefly before proceeding, causing a decrease in the speed, which is visible in Fig. 5. This behavior is consistent across teams and trajectories. UAV speed also has an important impact on signal measurement. Using the AFAR dataset, researchers can investigate the relationship between UAV speed, RSS fluctuations, and overall system performance, enabling more informed and adaptive algorithm development.